The table below show Business statistics I examination marks for a class of one hundred students .

Marks 10-19 (4 student )

20-29 ( 8 student )

30-39 ( a - student )

40-49 ( 22 student )

50-59 ( 48 student )

60-69 ( b - student )



The mean score of the students was 46.5

(i) Show that the values of a and b are 12 and 6 respectively. (5marks)

(ii) Estimate the median for the above sample. (3marks)

(iii) Determine the variance and the standard deviation (4 marks)

(iv) Given that 32% of the students failed the examination, what was the cut off points.

(3 marks)

(v)Estimate the mode for this data. (3marks)


1
Expert's answer
2021-10-19T13:18:02-0400

Marks frequency"(f)" class mark"(x)" "fx"

10-19 4 14.5 58

20-29 8 24.5 196

30-39 "a" 34.5 34.5a

40-49 22 44.5 979

50-59 48 54.5 2616

60-69 "b" 64.5 64.5b

Total frequency is given as, "\sum(f)=100".

Therefore,

"\sum(f)=4+8+a+22+48+b=100"

"82+a+b=100"

"a+b=18.........(i)"

Given that the mean "\bar{x}=46.5" and its formula is given by the formula'

"\bar{x}=\sum(fx)/\sum (f)"

Since we already have the quantity "\sum(f)", let us find "\sum(fx)" given by,

"\sum(fx)=(58+196+34.5a+979+2616+64.5b)=(3849+34.5a+64.5b)"

Now,

"\bar{x}=(3849+34.5a+64.5b)/100=46.5"

This implies that, "3849+34.5a+64.5b=4650"

Hence,

"34.5a+64.5b=801...............(ii)"

To find the values of unknowns "a" and "b," we solve for equations "(i)" and "(ii)" using substitution method.

In Equation "(i)" the value of "a" can be given as, "a=18-b". Substituting this value of "a" in equation "(ii)" gives,

"34.5(18-b)+64.5b=801"

"621-34.5b+64.5b=801"

"30b=180"

"b=6"

The value of "a" is found using the fact that "a=18-b" .

Therefore,

"a=18-6=12"

Thus, the values for "a" and "b" are 12 and 6 respectively as required.

b.

"Table 1"

Marks frequency"(f)" class mark"(x)" "fx" class boundary C.F

10-19 4 14.5 58 9.5-19.5 4

20-29 8 24.5 196 19.5-29.5 12

30-39 12 34.5 414 29.5-39.5 24

40-49 22 44.5 979 39.5-49.5 46

50-59 48 54.5 2616 49.5-59.5 94

60-69 6 64.5 387 59.5-69.5 100

C.F is the cumulative frequency.

To find the median, let us first determine the median class.

The median value is,

Median value="\sum(f)/2=100/2=50"

The median value is at the "50^{th}" position therefore the median class is, 50-59.

To find the median, we apply the formula below,

"median=l+((\sum(f)/2)-C.F)*c)/f" where,

"l" is the lower class boundary of the median class=49.5

"C.F" is the cumulative frequency of the class preceding the median class=46

"c" is the class width of the median class=10

"f" is the frequency of the median class.

Hence,

"median=49.5+((50-46)*10/48)=49.5+(40/48)=49.5+0.8333=50.333(3\space decimal \space places)"

c.

"Table 2"

Marks "f" "x" "fx" C.F "fx^2"

10-19 4 14.5 58 4 841

20-29 8 24.5 196 12 4802

30-39 12 34.5 414 24 14283

40-49 22 44.5 979 46 43565.5

50-59 48 54.5 2616 94 142572

60-69 6 64.5 387 100 24961.5

C.F is the cumulative frequency.

The variance is given by,

"variance=(1/(\sum(f)-1))*(\sum(fx^2)-(\sum(fx))^2/(\sum(f)))"

"variance=(1\/99)*(231025-(21622500\/100))"

"variance=(1/99)*(14800)=149.495(3\space decimal\space places)"

Standard deviation(SD) is given as,

"SD=\sqrt{variance}=\sqrt{149.495}=12.227(3\space decimal \space places)"

d.

To find the cut off points, we have to determine the "32^{nd}" percentile given as,

"P_{32}=l+(((32*\sum(f)/100))-C.F)*c/f" where,

"l" is the lower class boundary of the class containing "P_{32}"

"f" is the frequency of the class containing "P_{32}"

"c" is the width of the class containing "P_{32}"

"C.F" tis the cumulative frequency of the class preceding the class containing "P_{32}"

Let us find the class containing "P_{32}" given as,

"32*100\/100=32^{nd}" position. This position is in the class 40-49 as shown in "Table1".

We shall use "Table1" to tackle this question.

Now,

"P_{32}=39.5+(32-24)*10/22=39.5+3.6364=43.1364\approx43".

Therefore the cut off point is 43.

e.

To find the mode we shall use data in "Table1"

Mode is defined by the formula,

"mode=l+(f_m-f_1*c)\/(2*f_m-f_1-f_2)"

where,

"l" is the lower class boundary of the modal class

"c" is the class width of the modal class

"f_m" is the frequency of the modal class

"f_1" is the frequency of the class preceding the modal class

"f_2" is the frequency of the class succeeding the modal class

The modal class is the class with the highest frequency=48 thus, modal class is, 50-59

Therefore,

"mode=49.5+((48-22)*10)/(2*48-6-22)=49.5+(260/68)=49.5+3.8235=53.3235(4\space decimal \space places)"